# Kinetic Flat-Histogram Simulations of Non-Equilibrium Stochastic Processes with Continuous and Discontinuous Phase Transitions

### v1

@@ Abstract
As far as we know, there is no flat-histogram algorithm to sample the stationary distribution of non-equilibrium stochastic processes. The present work addresses this gap by introducing a generalization of the Wang-Landau algorithm, applied to non-equilibrium stochastic processes with local transitions. The main idea is to sample macroscopic states using a kinetic Monte Carlo algorithm to generate trial moves, which are accepted or rejected with a probability that depends inversely on the stationary distribution. The stationary distribution is refined through the simulation by a modification factor, leading to convergence toward the true stationary distribution. A visitation histogram is also accumulated, and the modification factor is updated when the histogram satisfies a flatness condition. The stationary distribution is obtained in the limit where the modification factor reaches a threshold value close to unity. To test the algorithm, we compare simulation results for several stochastic processes with theoretically known behavior. In addition, results from the kinetic flat-histogram algorithm are compared with standard exact stochastic simulations. We show that the kinetic flat-histogram algorithm can be applied to phase transitions in stochastic processes with bistability, which describe a wide range of phenomena such as epidemic spreading, population growth, chemical reactions, and consensus formation. With some adaptations, the kinetic flat-histogram algorithm can also be applied to stochastic models on lattices and complex networks.
@@ I Introduction
The Wang-Landau algorithm [1, 2, 3, 4, 5] is a stochastic sampling method that affords the estimation of the density of states $g(E)$ in statistical equilibrium models. The $g(E)$ is the relative degeneracy of a particular energy state and represents the statistical weight in the microcanonical ensemble.
While standard Monte Carlo methods, such as the Metropolis-Hastings algorithm, are based on importance sampling of the equilibrium canonical distribution at a particular temperature [5], the Wang-Landau algorithm is designed to sample the energy space uniformly. Also, it is related to reweighting techniques, while possessing the advantages of simplicity and straightforward implementation [5].
The main idea of the Wang-Landau algorithm is to perform a random walk in the space of energy states $E_{i}$ of equilibrium models. It starts with an initial guess for the density of states $g(E)=1$ and updates it in a non-Markovian way. Energy transitions are done according to the probability
@@
where $E_{i}$ and $E_{f}$ are the energy states before and after the trial, respectively. Following a transition, the density of states is updated by
@@
with $f$ being a modification factor which is initially set to a value greater than one, typically $f_{0}=e\sim 2.71828$. If a transition is accepted, we update the density of states at $E_{f}$; otherwise, we update the density of states at $E_{i}$.
@@
In addition, a histogram $H(E)$ is accumulated, which counts the number of visits to each energy state. The modification factor is reduced according to, for example, $f\to\sqrt{f}$ when the histogram satisfies a flatness criterion. The flatness criterion is typically defined such that the difference between any histogram bin and the average of the histogram does not exceed a given flatness threshold. The algorithm is stopped when the modification factor satisfies $1<f<f_{\text{final}}$, where $f_{\text{final}}$ is a modification factor threshold close to unity.
@@
A common problem with the Wang-Landau method is the error saturation [6]. One might naively expect that the modification factor convergence to unity would also lead to the convergence of $g(E)$. However, the deviation from the real density of states saturates, which can be alleviated by subtle modifications to the original algorithm [7, 6, 8]. Because of this error saturation, it is generally advised to compare Wang-Landau results with standard Monte Carlo averages. Progress is still being made in developing the Wang-Landau algorithm and extending its applications [9, 10, 11, 12, 13, 14].
@@
Also, one can see that the Wang-Landau algorithm does not satisfy the detailed balance condition in the early steps of the random walk. However, if the estimated density of states is sufficiently close to the real density of states, the detailed balance is recovered. The Wang-Landau algorithm allows direct calculation of the Helmholtz free energy and entropy of equilibrium systems, which are not directly accessible by standard Monte Carlo methods that sample the canonical ensemble [5].
@@
The Wang-Landau algorithm has been successfully applied to a wide range of equilibrium models, including spin systems [15, 4, 16, 17], lattice gases [18], protein folding [19], polymer models [20], and even quantum systems [9]. In addition, the Wang-Landau algorithm can be used to study systems with continuous degrees of freedom [10, 21] and with complex energy landscapes [2]. Also, the Wang-Landau algorithm enables direct evaluation of the latent heat of a bistable equilibrium system.
@@
However, as far as we know, there is no flat-histogram algorithm to sample the stationary distribution of non-equilibrium stochastic processes [22, 23, 24], some of which exhibit bistable behavior [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The present work addresses this gap by introducing a generalization of the Wang-Landau algorithm, applied to non-equilibrium stochastic processes with local transitions, that obey the following master equation:
$\frac{d}{dt}\mathcal{P}_{\boldsymbol{\sigma}}=\sum_{n=1}^{N}\left[w_{n}(\boldsymbol{\sigma}^{(n)})\mathcal{P}_{\boldsymbol{\sigma}^{(n)}}-w_{n}(\boldsymbol{\sigma})\mathcal{P}_{\boldsymbol{\sigma}}\right],$ (3)
@@
in the case of one degree of freedom. The algorithm can be generalized for stochastic processes with more than one degree of freedom. The master equation is a balance equation that gives $\mathcal{P}_{\boldsymbol{\sigma}}$, the probability distribution of the system states $\boldsymbol{\sigma}$, where each state is composed of $N$ local stochastic variables $\sigma_{n}$:
$\boldsymbol{\sigma}=\left(\sigma_{1},\sigma_{2},\ldots,\sigma_{n},\ldots,\sigma_{N}\right).$ (4)
@@
The master equation above describes Markovian dynamics with local transition rates $w_{n}(\boldsymbol{\sigma})$ among the system states. A local transition, i.e., a change of a local stochastic variable $\sigma_{n}\to\sigma^{\prime}_{n}$ with rate $w_{n}(\boldsymbol{\sigma})$, can produce a new system state $\boldsymbol{\sigma}^{(n)}$.
@@
$\frac{d\sigma}{dt}=-2\langle\sigma_{n}w_{n}(\boldsymbol{\sigma})\rangle,$ (5)
obtained directly from Eq. (3), where $P(\sigma)$ is related to the stationary system distribution by
$P(\sigma)=\sum_{\langle\boldsymbol{\sigma}\rangle=\sigma}^{\Omega}\mathcal{P}_{\boldsymbol{\sigma}},$ (6)
with $\Omega$ being the total number of states. As already stated, the method can be generalized for more than one degree of freedom. We consider cases where the degrees of freedom are bounded, and the dynamics can be entirely described by only one observable.
A potential application is the identification of discontinuous non-equilibrium phase transitions, i.e., non-equilibrium systems that display bistability. We present simulation results for some well-known bistable systems. A bistable non-equilibrium system can exhibit an exponential decay of the main observable at the threshold [36]. However, the kinetic flat-histogram algorithm can be used to directly analyze the stationary behavior.
The foremost distinctive feature of the kinetic flat-histogram algorithm is its extension of the Wang-Landau algorithm to systems that are not described by an energy landscape. In this approach, simulations are performed separately for each value of the control parameter (in equilibrium systems, this parameter is generally the temperature), which places the kinetic flat-histogram algorithm closer to multicanonical Monte Carlo algorithms [5]. Therefore, the obtained stationary distribution is $P(\sigma,\lambda)$, where $\lambda$ is the control parameter.
We organize this paper as follows. In Sec. II, we present the kinetic flat-histogram algorithm. In Sec. III, we show results of the algorithm for some mean-field stochastic processes with continuous phase transitions. In Sec. IV, we report findings for some bistable non-equilibrium stochastic processes. Finally, we provide our concluding remarks in Sec. V.
@@ II Kinetic Flat-Histogram Algorithm
In this section, we present the kinetic flat-histogram algorithm. In principle, the process should be Markovian with local updates, and the stochastic process should be ergodic. Some stochastic processes considered in this work do not fulfill this condition because they display absorbing states. The presence of absorbing states can lead to non-equilibrium phase transitions, where the most common universality class is the directed percolation [22, 23, 24]. Additionally, the stochastic process should obey the Perron-Frobenius theorem [37, 38], ensuring a well-defined stationary distribution.
One can address absorbing states by perturbing the non-equilibrium dynamics with reactivation [39], allowing the dynamics to continue when the system visits an absorbing state. Thus, the dynamics can be made ergodic to continue the simulation. The spontaneous particle insertion when the system visits the absorbing state is equivalent to an external field that scales as $1/N$ away from the threshold (in fact, the external field scales as $N^{-\sigma}$ at the threshold).
We regard well-studied stochastic processes in the mean-field regime, where exact results exist. By hypothesis, we can generate local transitions, referred to as trials, that update the main observable $\sigma_{i}\to\sigma_{f}$. Trials can be performed using standard Monte Carlo algorithms. For mean-field (one-site) stochastic processes, the exact stochastic simulation algorithm (SSA)[40, 41] is used. For lattice models and stochastic processes on graphs and networks[42], the kinetic Monte Carlo algorithm [5] can be employed.
In the case of one-site reactive processes with one degree of freedom [38], described by the reversible reaction
@@
where $A_{\ell}$ are one of the $q$ chemical species and $\nu_{i}^{\prime}$ and $\nu_{\ell}^{\prime\prime}$ are the respective stoichiometric coefficients, one can map Eq. (3) onto a birth-death process, which is a generalization of the Ehrenfest urn model with non-symmetric transition probabilities:
@@
with $a_{\sigma}$ and $b_{\sigma}$ being particular cases of the general transitions $w_{n}(\boldsymbol{\sigma})$.
Eq. (8), in the continuous limit, can be expressed as the following Fokker-Planck equation for the distribution of the main observable [38]:
$\frac{\partial}{\partial t}P(\sigma,t)$ $=$ $-\frac{\partial}{\partial\sigma}\left\{\left[a(\sigma)-b(\sigma)\right]P(\sigma,t)\right\}$ (9)
$+\frac{1}{2N}\frac{\partial^{2}}{\partial\sigma^{2}}\left\{\left[a(\sigma)+b(\sigma)\right]P(\sigma,t)\right\},$
where $P(\sigma,t)$ is the continuous limit of $P_{\sigma}(t)$ and $N$ is the system size. For one-site reactive processes, $a(\sigma)$ and $b(\sigma)$ in Eq. (9) can be calculated from
@@
where $x_{\ell}=N_{\ell}/N$ are the concentrations of the chemical species $A_{\ell}$, given as functions of the observable $\sigma$.
The kinetic flat-histogram algorithm samples the stationary solution of Eq. (8). The stationary distribution also satisfies Eq. (9) in the continuous limit, defined as the asymptotic $N\to\infty$ limit while the concentrations are still finite.
From Eq. (9), we can obtain the Langevin equation:
where $\left\langle\zeta(\sigma,t)\right\rangle=0$, and $\left\langle\zeta(\sigma,t)\zeta(\sigma,t^{\prime})\right\rangle=\delta(t-t^{\prime})$. The time-evolution equation for the main observable $\sigma$ can be obtained by simply averaging the Langevin equation.
Generalization to more than one degree of freedom (more reactions) and external sources is straightforward [38]. The results given in Eq. (10), Eq. (11), and their generalized versions for more than one degree of freedom are applied to the one-site stochastic processes of Secs. III and IV to obtain the time-evolution equations.
In general, the transition rates are functions of an external control parameter $\lambda$ that drives a phase transition, where the order parameter is the average of the main observable. For each value of $\lambda$, we perform a distinct simulation that yields a unique stationary distribution $P(\sigma,\lambda)$. Given the trials $\sigma_{i}\to\sigma_{f}$, the essence of the kinetic flat-histogram algorithm is as follows:
@@
    Initialize the system with random occupations, i.e., a random state. The algorithm updates a modification factor $f$, which starts with the value $\ln f_{0}=1$. Also, initialize two arrays: $\ln P(\sigma,\lambda)=0$ and $H(\sigma,\lambda)=0$, where we store the logarithm of the stationary distribution and the normalized visitation histogram for all possible values of the main observable $\sigma$, respectively. At the end of the simulation, $P(\sigma,\lambda)$ will converge to the true stationary distribution. Comparing expected values from the stationary distribution with time series averages from standard methods is required to test convergence;
@@
    Begin updating the main observable. Accept trials with the following probability:
@@
[…]
    respectively, where $\sigma=\sigma_{f}$ in the case of an accepted trial, and $\sigma=\sigma_{i}$ in the case of a rejected one. Also update the system state $\boldsymbol{\sigma}\to\boldsymbol{\sigma}^{(n)}$ and the value of the main observable $\sigma_{i}\to\sigma_{f}$ for an accepted trial. A Monte Carlo step is then defined as $N$ local updates;
@@
    After a number of Monte Carlo steps (we chose $10^{4}$ steps for all results), test a flatness condition for the histogram. The flatness condition is that all histogram bins should satisfy $0.95\langle H\rangle\leq H(E_{i})\leq 1.05\langle H\rangle$ where $\langle H\rangle$ is the histogram average. If the flatness condition is satisfied, update the modification factor as
@@
    End the simulation when $0<\ln f<\ln f_{\text{final}}$. In most of our results, we chose $\ln f_{\text{final}}=10^{-7}$.
We make some observations. First, one can implement dynamic arrays for the stationary distribution and the visitation histogram, growing the arrays according to the visitation. Second, we chose a “heat-bath” probability in Eq. (12) instead of a “Metropolis-Hastings” one as in the original Wang-Landau algorithm, given in Eq. (1).
After the simulation, we can calculate the following expected value $\varsigma(\lambda)$ for the main observable $\sigma$ from the converged stationary distribution $P(\sigma,\lambda)$:
@@
In addition, the kinetic flat-histogram algorithm allows direct calculation of the Shannon entropy, defined as
@@
In equilibrium systems that obey detailed balance, the Shannon entropy is the same as the Boltzmann entropy.
In the next sections, we show results for the algorithm applied to some well-known stochastic processes in the mean-field regime.

### v2

@@ Abstract
To our knowledge, there is no flat-histogram algorithm to sample the stationary distribution of non-equilibrium stochastic processes. The present work addresses this gap by introducing a generalization of the Wang-Landau algorithm, applied to non-equilibrium stochastic processes with local transitions. The main idea is to sample macroscopic states using a kinetic Monte Carlo algorithm to generate trial moves, which are accepted or rejected with a probability that depends inversely on the stationary distribution. The stationary distribution is refined through the simulation by a modification factor. A visitation histogram is also accumulated, and the modification factor is updated when the histogram satisfies a flatness condition. An estimate of the stationary distribution is obtained in the limit where the modification factor reaches a threshold value close to unity. To test the algorithm, we compare simulation results for several stochastic processes with theoretically known behavior. In addition, results from the kinetic flat-histogram algorithm are compared with standard exact stochastic simulations. We show that the kinetic flat-histogram algorithm can be applied to phase transitions in stochastic processes with bistability, which describe a wide range of phenomena such as epidemic spreading, population growth, chemical reactions, and consensus formation. With some adaptations, the kinetic flat-histogram algorithm can also be applied to stochastic models on lattices and complex networks.
@@ I Introduction
The Wang-Landau algorithm [83, 84, 46, 73, 45] is a stochastic sampling method that enables the estimation of the density of states $g(E)$ in statistical equilibrium models. Here, $g(E)$ denotes the relative degeneracy of a particular energy state and represents its statistical weight in the microcanonical ensemble.
Whereas standard Monte Carlo methods, such as the Metropolis-Hastings algorithm, rely on importance sampling of the equilibrium canonical distribution at a fixed temperature [45], the Wang-Landau algorithm is designed to sample the energy space uniformly. It is also related to reweighting techniques, while offering the advantages of simplicity and straightforward implementation [45].
The Wang-Landau algorithm performs a random walk in the space of energy states $E_{i}$ of equilibrium models. It begins with an initial guess for the density of states, $g(E)=1$, and updates it in a non-Markovian manner. Energy transitions are attempted with probability
@@
where $E_{i}$ and $E_{f}$ are the energy states before and after the trial, respectively. After each transition, the density of states is updated as
@@
where $f$ is a modification factor, initially set to a value greater than one, typically $f_{0}=e\approx 2.71828$. If a transition is accepted, the density of states is updated at $E_{f}$; otherwise, it is updated at $E_{i}$.
@@
Additionally, a histogram $H(E)$ is accumulated to count the number of visits to each energy state. The modification factor is reduced, for example, by $f\to\sqrt{f}$ whenever the histogram satisfies a flatness criterion. This criterion is typically defined such that the difference between any histogram bin and the average histogram value does not exceed a specified threshold. The algorithm terminates when the modification factor satisfies $1<f<f_{\text{final}}$, where $f_{\text{final}}$ is a threshold value close to unity.
@@
A common issue with the Wang-Landau method is error saturation [7]. One might expect that as the modification factor converges to unity, $g(E)$ would also converge. However, the deviation from the true density of states saturates, which can be mitigated by subtle modifications to the original algorithm [6, 7, 88]. Due to this error saturation, it is generally recommended to compare Wang-Landau results with standard Monte Carlo averages. Ongoing research continues to improve the Wang-Landau algorithm and expand its applications [72, 87, 58, 80, 12, 22, 47]. Another example of a flat-histogram algorithm is the multicanonical Monte Carlo method [48, 49, 50, 51], which has been shown not to suffer from error saturation.
@@
Additionally, it is important to note that the Wang-Landau algorithm does not satisfy the detailed balance condition during the initial stages of the random walk [83, 46]. However, as the estimated density of states approaches the true density of states, detailed balance is restored for the stationary density of states. The Wang-Landau algorithm enables the direct calculation of the Helmholtz free energy and entropy in equilibrium systems, quantities that are not readily accessible through standard Monte Carlo methods sampling the canonical ensemble [45].
@@
The Wang-Landau algorithm has been successfully applied to a wide range of equilibrium models, including spin systems [69, 73, 40, 41], lattice gases [37], protein folding [66], polymer models [14], and quantum systems [72]. Furthermore, it can be used to study systems with continuous degrees of freedom [87, 70] and complex energy landscapes [84, 48, 49, 50, 51]. The Wang-Landau algorithm also enables the direct evaluation of the latent heat in bistable equilibrium systems.
@@
However, to our knowledge, there is no flat-histogram algorithm for sampling the stationary distribution of non-equilibrium stochastic processes [63, 36, 34], some of which exhibit bistable behavior [68, 31, 32, 82, 89, 81, 19, 18, 23, 35, 79]. This work addresses this gap by introducing a generalization of the Wang-Landau algorithm, applied to non-equilibrium stochastic processes with local transitions that obey the following master equation
$\frac{d}{dt}\mathcal{P}(\boldsymbol{\sigma},t)=\sum_{n=1}^{N}$ $\left[w_{n}\!\left(\boldsymbol{\sigma}^{(n)}\to\boldsymbol{\sigma}\right)\,\mathcal{P}(\boldsymbol{\sigma}^{(n)},t)\right$. (3)
$-\left.w_{n}\!\left(\boldsymbol{\sigma}\to\boldsymbol{\sigma}^{(n)}\right)\,\mathcal{P}(\boldsymbol{\sigma},t)\right].$
@@
for the case of one degree of freedom. Generalization to stochastic processes with more than one degree of freedom is straightforward. The master equation is a balance equation that gives $\mathcal{P}_{\boldsymbol{\sigma}}$, the probability distribution of the system states $\boldsymbol{\sigma}$ which are given as
$\boldsymbol{\sigma}=\left(\sigma_{1},\sigma_{2},\ldots,\sigma_{n},\ldots,\sigma_{N}\right),$ (4)
@@
where each state is composed of $N$ local stochastic variables $\sigma_{i}$ ($i=1,2,\ldots,N$). Moreover, the master equation above describes Markovian dynamics with local transition rates $w_{n}(\boldsymbol{\sigma})$ among the system states. A local transition, i.e., a change of a local stochastic variable $\sigma_{n}\to\sigma^{\prime}_{n}$ with rate $w_{n}(\boldsymbol{\sigma})$, yields a new system state $\boldsymbol{\sigma}^{(n)}$ written as
$\boldsymbol{\sigma}^{(n)}=\left(\sigma_{1},\sigma_{2},\ldots,\sigma^{\prime}_{n},\ldots,\sigma_{N}\right).$ (5)
@@
$\frac{d\sigma}{dt}=-2\left\langle\sigma_{n}w_{n}\!\left(\boldsymbol{\sigma}\to\boldsymbol{\sigma}^{(n)}\right)\right\rangle,$ (6)
obtained directly from Eq. (3) for a homogeneous system, where $P(\sigma)$ is related to the stationary system distribution $\mathcal{P}(\boldsymbol{\sigma})$ by
$P(\sigma)=\sum_{\langle\boldsymbol{\sigma}\rangle=\sigma}^{\Omega}\mathcal{P}(\boldsymbol{\sigma}),$ (7)
with $\Omega$ being the total number of states. We then introduce weights in the random walk transitions to encourage visits to less frequently visited states, and the stationary distribution of the observable $\sigma$ is obtained from these weights when the visitation histogram becomes flat. Generalization to non-homogeneous systems is straightforward. We consider cases where the degrees of freedom are bounded and the dynamics can be entirely described by a single observable.
The concept of equilibrating transition rates is similar to that in Ref. [52]. However, the kinetic flat-histogram algorithm does not sample the Gibbs free energy, nor do the rates depend on the energy. Additionally, we maintain a histogram of the observable $\sigma$. Some of the tested models are described solely by master equations and do not possess an associated Hamiltonian.
A potential application is the identification of discontinuous non-equilibrium phase transitions, i.e., non-equilibrium systems that exhibit bistability. We present simulation results for several well-known bistable systems. A bistable non-equilibrium system can display an exponential decay of the main observable at the threshold [65]. However, the kinetic flat-histogram algorithm enables direct analysis of the stationary behavior.
The most distinctive feature of the kinetic flat-histogram algorithm is its extension of the Wang-Landau algorithm to systems not described by an energy landscape. In this approach, simulations are performed separately for each value of the control parameter (in equilibrium systems, this parameter is typically the temperature), which makes the kinetic flat-histogram algorithm more closely related to multicanonical Monte Carlo algorithms [45]. Consequently, the stationary distribution obtained is $P(\sigma,\lambda)$, where $\lambda$ is the control parameter.
We organize this paper as follows. In Sec. II, we present the kinetic flat-histogram algorithm. In Sec. III, we show results for several mean-field stochastic processes with continuous phase transitions. In Sec. IV, we report findings for some bistable non-equilibrium stochastic processes. Finally, we provide our concluding remarks in Sec. V.
@@ II Kinetic Flat-Histogram Algorithm
In this section, we present the kinetic flat-histogram algorithm. In principle, the process should be Markovian with local updates, and the stochastic process should be ergodic. Some stochastic processes considered in this work do not fulfill this condition because they display absorbing states. The presence of absorbing states can lead to non-equilibrium phase transitions, with the most common universality class being directed percolation [63, 36, 34]. Additionally, the stochastic process should obey the Perron-Frobenius theorem [74, 71], ensuring a well-defined stationary distribution.
Absorbing states can be addressed by perturbing the non-equilibrium dynamics with reactivation [57], which prevents the dynamics from being trapped. This approach ensures ergodicity and allows the simulation to continue. The main effect is that the stationary distribution is no longer trivial, since any dynamics would otherwise be trapped in the absorbing state for finite system sizes. Spontaneous particle insertion when the system visits the absorbing state is equivalent to an external field that scales as $1/N$ away from the threshold (in fact, the external field scales as $N^{-\sigma}$ at the threshold).
We consider well-studied stochastic processes in the mean-field regime, where exact results exist. By hypothesis, we can generate local transitions, referred to as trials, that update the main observable $\sigma_{i}\to\sigma_{f}$. Trials can be performed using standard Monte Carlo algorithms. For mean-field (one-site) stochastic processes, the exact stochastic simulation algorithm (SSA) [30, 29] is used. For lattice models and stochastic processes on graphs and networks [61], the kinetic Monte Carlo algorithm [45] can be employed.
In the following sections, we discuss several reactive processes as prototypes of non-equilibrium stochastic systems with phase transitions. For one-site reactive processes with a single degree of freedom [71], described by the reversible reaction
@@
where $A_{\ell}$ are the $q$ chemical species and $\nu_{\ell}^{\prime}$, $\nu_{\ell}^{\prime\prime}$ are the respective stoichiometric coefficients, Eq. (3) can be rewritten as
@@
where $a_{\sigma}$ and $b_{\sigma}$ are particular cases of the general transition rates $w_{n}$. Eq. (9) describes a standard birth-death process [71].
From Eq. (9), in the continuous limit, we can obtain the following Langevin equation for the main observable
where $\left\langle\zeta(\sigma,t)\right\rangle=0$ and $\left\langle\zeta(\sigma,t)\zeta(\sigma,t^{\prime})\right\rangle=\delta(t-t^{\prime})$. For one-site reactive processes, $a(\sigma)$ and $b(\sigma)$ in Eq. (10) can be calculated as [71]
@@
where $x_{\ell}=N_{\ell}/N$ are the concentrations of the chemical species $A_{\ell}$, given as functions of the observable $\sigma$. The time-evolution equation for the main observable $\sigma$ can be obtained by averaging the Langevin equation. Thus, the stochastic process admits two equivalent descriptions: the master equation and the Langevin equation, both of which can be addressed using the kinetic flat-histogram algorithm.
Generalization to more than one degree of freedom (i.e., additional reactions) and external sources is straightforward [71]. The result presented in Eq. (10), and their generalized forms for multiple degrees of freedom are applied to the one-site stochastic processes discussed in Secs. III and IV to obtain the corresponding time-evolution equations.
In general, the transition rates are functions of an external control parameter $\lambda$ that drives a phase transition, where the order parameter is the average of the main observable. For each value of $\lambda$, we perform a distinct simulation that yields a unique stationary distribution $P(\sigma,\lambda)$, which is analogous to simulate the original Wang-Landau algorithm for various values of the magnetic field or the exchange coupling constant in a magnetic system. Given the trials $\sigma_{i}\to\sigma_{f}$, the essence of the kinetic flat-histogram algorithm is as follows:
@@
    Initialize the system in a random state. Set the modification factor $f$ with $\ln f_{0}=1$. Also, initialize two arrays: $\ln P(\sigma,\lambda)=0$ and $H(\sigma,\lambda)=0$, which store the logarithm of the stationary distribution and the normalized visitation histogram for all possible values of the main observable $\sigma$, respectively. The arrays may be grown dynamically; in this case, a third array containing the visited $\sigma$ bins is added. Arrays should be resized and reordered when a new $\sigma$ is visited. At the end of the simulation, $P(\sigma,\lambda)$ will converge to the asymptotic stationary distribution. Comparing expected values from the stationary distribution with time-series averages from standard methods is recommended to test consistency;
@@
    Begin updating the main observable. Accept trials with the probability
@@
    $\left\{\begin{aligned} &\ln P(\sigma,\lambda)\to\ln P(\sigma,\lambda)+\ln f,\\ &H(\sigma,\lambda)\to H(\sigma,\lambda)+1,\end{aligned}\right$. (13)
    respectively, where $\sigma=\sigma_{f}$ if the trial is accepted, and $\sigma=\sigma_{i}$ if rejected. Also update the system state $\boldsymbol{\sigma}\to\boldsymbol{\sigma}^{(n)}$ and the value of the main observable $\sigma_{i}\to\sigma_{f}$ for accepted trials. A Monte Carlo step is defined as $N$ local updates;
@@
    After a number of Monte Carlo steps (we used $10^{4}$ steps for all results), test the flatness condition for the histogram. The flatness condition requires that all histogram bins satisfy $0.95\langle H\rangle\leq H(E_{i})\leq 1.05\langle H\rangle$, where $\langle H\rangle$ is the histogram average. If the flatness condition is satisfied, update the modification factor as
@@
    End the simulation when $0<\ln f<\ln f_{\text{final}}$. In most results, we chose $\ln f_{\text{final}}=10^{-7}$.
After the simulation, we can calculate the expected value $\varsigma(\lambda)$ of the main observable $\sigma$ from the converged stationary distribution $P(\sigma,\lambda)$ as
@@
Additionally, the kinetic flat-histogram algorithm enables the direct calculation of the Shannon entropy, defined as
@@
For equilibrium systems that obey detailed balance, the Shannon entropy coincides with the Boltzmann entropy.
In the following sections, we present results for the algorithm applied to several well-known stochastic processes.